OpenStudy (anonymous):

How would you find the exact values of the six trigonometric functions of the given angle 2pi/3?

OpenStudy (anonymous):

locate \[\frac{2\pi}{3}\] on the unit circle, first coordinate is cosine, second is sine, etc. a good unit circle is here http://tutorial.math.lamar.edu/cheat_table.aspx go to last page of trig cheat sheet

OpenStudy (anonymous):



OpenStudy (anonymous):

your in quadrant 2 so sine is positive! not cos sorry. Use the special triangle and then inverse your ratios, sin=opp/hyp, cos=adj/hyp, tan=opp,adj